OpenStudy (anonymous):

solve equation 3x + 3 = 5x − 1

OpenStudy (anonymous):

ok i got this XD

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

Step 1: Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 3*x+3-(5*x-1)=0

OpenStudy (anonymous):

sam please dnt give direct answers how are they soppose to learn but ur correct

OpenStudy (anonymous):

can you show me how you got your answer

OpenStudy (anonymous):

ok yes i can

OpenStudy (anonymous):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 3*x+3-(5*x-1)=0

OpenStudy (anonymous):

3x+3=5x-1 add 1 to both sides leaving 3x+4=5x subtract 3x from both sides leaving 4=2x divide 2 from both sides

OpenStudy (anonymous):

Pull out like factors : -2x + 4 = -2 • (x - 2)

OpenStudy (anonymous):

then Solve -2*(x-2) = 0

OpenStudy (anonymous):

Solve : -2 = 0

OpenStudy (anonymous):

so you add the 1 to 3x and 3

OpenStudy (anonymous):

No just 3

OpenStudy (anonymous):

This equation has no solution.

OpenStudy (anonymous):

It dose have one how you are doing its it dosnt

OpenStudy (anonymous):

A a non-zero constant never equals zero.

OpenStudy (anonymous):

Solve : x-2 = 0 Add 2 to both sides of the equation : x = 2

OpenStudy (anonymous):